### v1

# Topological phases in discrete stochastic systems

Abstract
Topological invariants have proved useful for analyzing emergent function as they characterize a property of the entire system, and are insensitive to local details, disorder, and noise. They support edge states, which reduce the system response to a lower dimensional space and offer a mechanism for the emergence of global cycles within a large phase space. Topological invariants have been heavily studied in quantum electronic systems and been observed in other classical platforms such as mechanical lattices. However, this framework largely describes equilibrium systems within an ordered crystalline lattice, whereas biological systems are often strongly non-equilibrium with stochastic components. We review recent developments in topological states in discrete stochastic models in 1d and 2d systems, and initial progress in identifying testable signature of topological states in molecular systems and ecology. These models further provide simple principles for targeted dynamics in synthetic systems or in the engineering of reconfigurable materials. Lastly, we describe novel theoretical properties of these systems such as the necessity for non-Hermiticity in permitting edge states, as well as new analytical tools to reveal these properties. The emerging developments shed light on fundamental principles for non-equilibrium systems and topological protection enabling robust biological function.
I Introduction
Understanding how the underlying components of proteins, RNAs, metabolites, and other biologically relevant molecules give rise to biological function would enable precise targeting of interventions to enhance health. However, we still do not have a good theory for how structure leads to function in these complex and noisy systems. As just one example, despite successful sequencing of the human genome, it remains challenging to predict the behavior of resulting proteins and macromolecules due to the large space of possible configurations and reactions causing transitions between them [1, 2]. This large phase space of possible configurations also renders unfeasible exhaustive searches using other approaches like experiment or numerical simulation [3], underscoring the need for simple conceptual methods to provide insight [4, 5, 6].
Drawing on recent theoretical advances in topology, molecular biology, and non-equilibrium statistical physics, the study of topological states in discrete stochastic models provides a theoretical approach to predict robust and dynamic function in living systems. The origins of topological classification stem from random matrix theory [7, 8, 9]. These establish the necessary conditions for the existence of models with boundary responses that are robust to generic or random perturbations [10, 11]. In practice, such models are composed of repeated and common motifs (unit cells). This allows for the flexible design of networks that exhibit behavior on a submanifold of the larger system, without requiring fine-tuning or over-parametrization for this specialized behavior on the submanifold.
Such topological invariants have proved useful for analyzing emergent function as they characterize a property of the entire system [10, 12, 13], i.e., whether distinct chemical or mechanical configurations have a global pattern of transition rates between them that supports edge states. Edge states reduce the system response to a lower dimensional space and offer a mechanism for the emergence of global cycles within a large configuration space. Powerfully, this edge response is insensitive to local details, disorder, and noise, and can describe robust biological oscillations or growth cycles. Such theories could shed light on the fundamental question of why biological function is so robust, e.g. during development or when maintaining stable dynamics over long times, even in the presence of stochasticity or changing external conditions and stimuli.
Topological approaches have gained popularity in describing biological systems such as knots in DNA [14] and defects in tissues [15, 16]. Among these approaches, the framework of topological invariants can identify the global state of the system and its associated response, for large classes of different systems [10, 17, 18, 19], see Fig. 1. This framework has led to radical discoveries in other systems, most notably in quantum electrons [10, 12, 17, 20, 21, 22, 23], such as the discovery of dissipationless edge states that enabled precise measurements enabling new standards in metrology [18]. This framework has since been extended to various classical systems, from mechanical lattices [24, 25, 26] to photonics [27, 28, 29, 30], acoustics [31, 32], and active matter [33, 34, 35, 36, 37]. Most of these systems show high spatial order, or even an actual lattice structure in space, which is typically absent in biological systems that are very heterogeneous.
Recently, new mappings of topological tools for non-equilibrium stochastic systems have been demonstrated, and shown to yield localization [35, 34] and currents [33] on the boundaries of an abstract reaction or configuration space. They have been used to model dimensional reduction in the circadian rhythm [33, 40], sensory adaptation and kinetic proofreading [35], and population dynamics [36, 37]. Such models also provide sufficient building blocks for the design of synthetic biological or robotic systems [33].
In this review, we discuss different models that have been proposed in recent years, along with a discussion of attractive properties of these models as well as the formalism used to describe them. In turn, this burgeoning field of study has also prompted the development of new theoretical and computational approaches, as well as the discovery of new physical phenomena. We survey these and discuss the relevant mathematical approaches. We close with a discussion of potential applications to biological and non-biological systems, which herald exciting future directions for the field.
The review is organized as follows. In Section II, we give a brief overview of the development of topological phases initially for quantum matter, and later for classical systems in continuous, real space. In Section III, we describe recent discoveries of topological phases in discrete classical systems, where the dynamics take place in an abstract configuration space. These include stochastic dynamics in biochemical systems, and nonlinear deterministic dynamics in ecological systems. In the remainder of the review, we expand on the novel properties (Section IV), applications (Section V), and theoretical insights and tools (Section VI) that arise in stochastic systems in particular, although much of what we discuss also applies to population dynamics models.
II Background
II.1 Topological phases in quantum systems
Topology is concerned with the properties of objects that remain invariant under continuous deformations. While this can sound abstract, the study of topology has made powerful contributions to various physical problems, by distinguishing broad classes of materials and their resulting physical response through analysis of their fundamental properties. Its contribution to the field of quantum electronic systems has been striking, predicting which classes of materials exhibit currents or polarization localized to just the edges of the system [10, 12, 13, 19]. These edge currents are insensitive to local information such as impurities or defects, hence displaying such precise values that new standards for the measurement of fundamental constants have been established [18]. Due to this robust protection from disorder or noise, topological states have also been touted as candidates for fault-tolerant quantum computing [41].
The framework for topological states has been made rigorous thanks to the identification of topological invariants that govern the system. This is due to the celebrated bulk-boundary correspondence, where properties of the bulk (system interior) predict a specific and quantitative response on the boundary [10, 12, 19, 42]. Indeed, materials can be characterized just based on their dimensions and symmetry group [12], through systematic classifications which have been extended to interacting systems [43, 44]. Despite these elegant and fundamental theoretical predictions, the accessibility and detection of topological states can be challenging in practice. Very clean, cold, and ordered samples are often required to observe these delicate quantum properties [10, 13].
II.2 Topological phases in classical systems
Topological approaches in biology have mostly been based on identifying winding structures in biological data, including topological defects in tissues [16, 45, 46, 15, 47] or knots in DNA strands [14], and more abstractly in any high-dimensional dataset through methods such as persistent homology [48, 49, 50, 51, 52, 53]. Other approaches have linked the geometric phase of a system with its dynamics or response, in the context of locomoting or rotating animals [54, 55, 56] or stochastic chemical reactions [57].
The last decade has seen diverse proposals for realizations of new phases of matter characterized by topological invariants in classical platforms. These display edge localization and currents analogous to those in quantum systems, but can be probed in much less extreme conditions. Examples include mechanical lattices [24, 25, 26], photonics [27, 28, 29, 30], acoustics [31, 32], and active matter [33, 34, 35].
In all these examples of topological phases in classical systems, the dynamics of interest occur in actual physical space (e.g. mechanical lattices, optical waveguides, microfluidic devices…) and are described by continuum mechanics, typically in a deterministic setup. In contrast, the dynamics of interest in biological systems often do not take place in physical space, but rather in an abstract and typically discrete configuration space representing concentrations of chemicals, conformational states of proteins, or population numbers of individuals. These dynamics are often stochastic and driven by chemical reactions or birth-death processes. Thus, a new approach is needed to identify and describe topological phases in these systems, which we describe in the following.
III Topological phases in discrete classical systems
Repeated processes or motifs in configuration space, describing a basic reaction such as the conversion of biomolecules from one state to another, or the birth or death of individuals in a population, can be linked together to form an ordered network or lattice that allows for topological analysis. Most importantly, when many such motifs are linked together, it becomes possible to distinguish between the “bulk” of the lattice and its “edges”, the latter arising e.g. from a limiting supply of building blocks or other constraints in the reaction dynamics, that limit the extent of the configuration space. This mapping has been successfully exploited in recent years to allow for the rigorous analysis of topological invariants in stochastic and biological systems, crucially extending models for equilibrium systems to a range of phenomena that is strongly non-equilibrium. They provide a description for the emergence of robust global responses from many basic building blocks (see Fig. 2, top row).
A common example for a building block in the biochemical context is that of futile cycles [38, 39]; see Fig. 2, top left. Futile cycles consist of a series of chemical reactions that bring the molecule back to its starting point, a non-equilibrium process which dissipates energy and thus which requires the input of chemical fuel, typically provided by ATP or GTP in the biological context. Despite seeming wasteful (or futile, as their name suggests), such cycles can be found in protein synthesis, muscular contraction, metabolism, and sensory systems – where their purpose across all these disparate systems remains unclear [39]. Intriguingly, the stochastic topological models that we describe below show that systems with futile cycles as building blocks tend to display topological states [35, 34, 33] . This suggests a fascinating route by which seemingly wasteful non-equilibrium motifs in biology can collectively foster robust global function.
A similar role to that of futile cycles has been found to be played by rock-paper-scissors motifs in ecological networks describing population dynamics [36, 37]; see Fig. 2, top right. These rock-paper-scissors motifs represent a non-transitive relation among three competing species: rock outcompetes scissors, which outcompetes paper, which outcompetes rock [58, 59]. Non-transitive competition is widely prevalent in animals, plants, and microbes. As we describe below, when many of these rock-paper-scissors cycles are linked together, representing cases where a species plays a role in more than one such cycle within a larger ecosystem, topological states emerge at the scale of the whole ecosystem.
III.1 Stochastic systems
The stochastic dynamics of biochemical systems are typically described by continuous time Markov processes on a discrete state space, through a master equation of the form , where is a vector containing the probabilities of being in state (normalized as ), and is a transition matrix where is the transition rate (probability per unit time) of the transition [60]. To enforce conservation of probability, the transition matrix must satisfy for and . These properties (together with the requirement that the directed graph corresponding to the network of transitions is strongly connected, so that there are no sinks or disconnected regions) guarantee that the transition matrix has one zero eigenvalue, corresponding to the steady state which satisfies , while all other eigenvalues have negative real part, representing exponentially decaying modes (which may be oscillatory or non-oscillatory depending on whether the imaginary part is nonzero or zero, respectively). This formalism is widely used to model conformational changes of proteins and protein complexes [61], the operation of molecular motors [62, 63] and stochastic swimmers [64, 65], and the full stochastic dynamics of chemical reactions in finite systems [66].
III.1.1 1D systems
The first demonstration of topological states in stochastic systems was given for one-dimensional (1D) lattices in Ref. 35 and followed-up on shortly after in Ref. 34. A basic depiction of the model and results is shown in Fig. 2(a). The basic lattice has a “ladder” shape, with unit cells containing two sites (top and bottom) that are replicated laterally along the horizontal direction to form a 1D bulk. Considering the transitions (both ways) among internal sites of the unit cell, and among top sites and bottom sites in neighboring cell, a given bulk is defined by six transition rates. The authors considered what occurs at the interface between two different lattices (defined by two different bulks, i.e. two different sets of transition rates) when the two are linked together.
The authors showed that, for any given bulk, one can define a topological index, described further below. Crucially, when the topological indexes of the two bulks that are brought into contact do not coincide, localization (i.e. accumulation of probability at steady state) is observe at the interface between the two.
In summary, these seminal works demonstrated that topological considerations based purely on bulk properties can be used to predict global effect, occurring at the edges of a lattice or at the interfaces between two connected lattices. However, because of the 1D nature of the systems, the topological phenomena that could be observed were limited to static localization.
III.1.2 2D systems
These restrictions were lifted by us in Ref. 33, which provided the first realization of topological states in two-dimensional (2D) stochastic systems. In 2D systems, besides static polarization, topologically-protected chiral currents along the edges of the system, analogous to those seen in the quantum Hall effect [19], become possible.
In the minimal model, see Fig. 2(b), the unit cell has four sites, labelled A–D, and is replicated along two directions forming a 2D lattice (note that a three-site model can also be defined [33], but is less amenable to analysis). The four sites of a unit cell are connected by four transitions forming an “internal” futile cycle. In the bulk of the system, four neighboring unit cells are also connected by four transitions forming an “external” futile cycle. The minimal model, with unidirectional transitions, thus has eight transition rates in total (which become sixteen transition rates if bidirectional transitions are allowed). However, one can further impose rotational symmetry and make all internal transitions equal to each other, and similarly for external transitions, so that only two transition rates (internal and external) remain and define the bulk behavior.
Importantly, when (i) internal and external cycles have opposite chirality and (ii) internal transitions are much slower than external ones, probability accumulates at the edges of the system, and a probability current is observed along these edges, with the same chirality as that of the internal cycles. Thus, global oscillations are observed at a system-spanning scale. Further analysis, described below, shows that this effect is topological and thus the edge currents are protected against disorder and deformations of the system boundary.
III.2 Population dynamics
Population dynamics models are typically described at the mean field level, with non-linear equations that describe the number, concentration, or fraction of individuals of different species in a well-mixed (spatially homogeneous) ecosystem. The simplest and most commonly studied description for the competition between different species is the Lotka-Volterra Equation [67]. For the particular case of a zero-sum game in which the total population is conserved, and a decrease in the population of one species is balanced by an increase in the population of another, the deterministic dynamics are described by the Antsymmetric Lotka-Volterra Equation (ALVE), which can be written as , where is the population of species , and is an antisymmetric interaction matrix with components that define the transfer of population between species and [67, 68]. The total population is conserved, and without loss of generality can be normalized to . The ALVE also appears in various contexts other than population dynamics, such as game theory, Bose-Einstein condensation, plasma physics, and chemical kinetics [68]. Note that, while bearing some similarities to the master equation described above, as both describe the dynamics of a conserved quantity that is transferred across discrete sites, the ALVE describes very different dynamics: first, it is nonlinear (as it describes a many-body interacting system at the mean field level) while the master equation is linear; second, the matrix is antisymmetric while has no such constraint, although it does have other constraints.
III.2.1 1D systems
As in the case of biochemical systems, the first demonstration of topological states in population dynamics was given for a 1D lattice in Ref. 36. The authors linked together many rock-paper-scissors cycles into a long chain, see Fig. 2(c). The bulk of the system is thus defined by the three interaction strengths that define the cycle. By construction, in each cycle there is always one ‘isolated’ species that only interacts internally with the other two species in the cycle, and two ‘connected’ species that also take part in the neighboring cycle, and thus interact with four species in total. The authors found that, depending on the skewness of the cycle, determined by the ratio of interaction strengths between the isolated and the two connected species, population would accumulate on either the left or the right end of the chain. Thus, similarly to the 1D biochemical systems, static polarization is observed at steady state. Importantly, polarization was robust to initial conditions and to small disorder in the system parameters, suggesting a topological origin of the transition. This was confirmed by a mapping of the interaction matrix to a Hermitian Hamiltonian that could be analyzed using standard tools of quantum topology, described further below. We note as well that the effect of deviations from ideal rock-paper-scissors interactions in the 1D chain was recently studied in Ref. 69.
III.2.2 2D and 3D systems
In Ref. 37, a 2D Kagome lattice of rock-paper-scissors cycles was studied, see Fig. 2(d). All interactions were chosen to be identical, so that the system is essentially parameter-free. The authors found that, when the system is initialized with additional population in one of the edge sites of the lattice, this additional population propagates along the edges of the lattice, with the same chirality as that of the rock-paper-scissors cycles. As in 2D biochemical systems, this edge current was seen to be topologically protected, and in particular to be robust to deformations of the shape of the boundaries. By mapping the interaction matrix to a Hermitian Hamilonian in a similar way as for the 1D chain, they could show that the edge current is induced by a nontrivial topological index (Chern number).
By stacking several of these 2D Kagome lattices on top of each other along a third () axis, a 3D lattice of rock-paper-scissors cycles was constructed in Ref. 70. Similarly to the 2D case, chiral propagation of population density was observed in the dynamics projected onto the plane, implying density propagation along the faces of the system. No significant dynamics were reported along the dimension. The chiral propagation was shown to be robust and to have a topological origin, through a similar mapping as for the 1D and 2D cases.
III.3 Formalism and mapping to quantum systems
Topological invariants in stochastic systems can be identified using results developed in quantum systems. In Ref. 35, a topological index was identified from the winding number in Fourier space of the phase associated to the determinant of the tilted current matrix associated to the bulk. In follow-up work, this index was further related to the topological index of the Hermitian Hamiltonian of an associated supersymmetric quantum system [34].
A more direct method that flexibly maps many systems described by the master equation to a quantum Hamiltonian was used in Ref. 33, although the resulting Hamiltonian in this case is typically non-Hermitian [71, 72]. Namely, the similarity of the master equation to the Schrödinger equation is exploited, as they are formally identical up to the constant . In quantum systems, topological invariants such as the Chern number are typically calculated from the eigenvectors of the Hamiltonian . When mapping a master equation to a corresponding Schrödinger equation, the eigenvectors of the transition matrix will be identical to those of , in the case of periodic boundary conditions. Hence, the same calculation that verifies if the quantum system is topological using its eigenvectors, can be performed for from a stochastic system. Using the bulk-boundary correspondence, a topological invariant that characterizes the bulk will predict a boundary response for the system. Further works have shown that this bulk-boundary correspondence has additional constraints in stochastic systems as compared to quantum ones [73, 74, 75], which is discussed below in Section VI.
A similar approach can be used in nonlinear models of population dynamics described by the ALVE. In particular, when the ALVE is linearized around a uniform steady state solution, with deviations , where is the number of species in the lattice, the equation can be written as , which has the same form as the Schrödinger equation with a Hamiltonian [37]. In this case, because the interaction matrix is antisymmetric, the Hamiltonian is Hermitian, and the usual tools of Hermitian quantum topology can be used [36, 37, 70].
IV Attractive features and useful properties
IV.1 Robustness
Topological states are known to be robust to deformations of the edge, and hence to impurities or disorder. The global response in such stochastic systems therefore inherits the useful property of topological protection that provides robustness to missing components or obstacles, see Fig. 3 [33, 40]. This behavior can shed light on how biological function remains so robust, e.g. during development or when maintaining stable dynamics over long times, even in the presence of stochasticity or changing external conditions and stimuli.
Numerical simulations have been conducted to test the robustness of these topological models to random noise or perturbations, which are common in biological systems. By systematically adding random fluctuations to transition rates, these simulations have shown that weak disorder does not destroy the topological phase [35, 40].
IV.2 Dimensional reduction of responses and dynamics
The reduction of the full system response to a much lower dimensional description has been observed in many complex biological systems, where the system dynamics or behavior is confined to a much smaller subset of the phase space [76, 77, 78]. This is exemplified in computational models of memory, that describe specific attractor states which represent persistent memories [79, 80]. Another example is that of long oscillations, such as the circadian rhythm, which are crucial for the regulation of many processes such as metabolism and replication [81, 82]. Confining the dynamics onto the lower dimensional boundary of a much larger network, topological models in stochastic systems offer a route to achieve this ubiquitous phenomenon.
As topological models demonstrate how to achieve dimensional reduction in a precise manner, stochastic versions provide the necessary and sufficient conditions under which biological networks, notoriously large and messy, can reduce their activity to just a submanifold of their full phase space. Moreover, because edge states decay exponentially away from the system edge with typically short decay lengths [33], they can persist even in small systems. Short decay lengths also make it possible for domains with distinct topological invariants to co-exist on the same network, allowing for localized states or currents at the boundary between subnetworks or subregions.
IV.3 Small motifs that can be scalable
Topological models typically consist of lattices, which provide a versatile description as they consist of minimal biochemical motifs describing just a few interlinked reactions (see Fig. 2). These underlying building blocks are repeated across the network and support the emergence of robust global responses. As the same motifs are repeated, the resulting responses are governed by just a few parameters. Meanwhile, the repeated motifs can scale up towards large networks and systems, providing a flexible design architecture.
V Applications to biology and synthetic systems
The discrete nature of molecules and proteins, which moreover often exist only in small copy number, naturally lends itself to discrete stochastic descriptions. In these systems, the stochastic nature of discrete events can yield new physics, such as extinction events in gene regulation [83]. Topological modeling of this new physics can provide insights about robustness in chemical and biological systems, that can be further extended to synthetic systems.
V.1 Molecular biology
Non-equilibrium motifs abound in many different biochemical systems, making the formalism described here applicable to macromolecular or cellular dynamics. It is thus desirable to investigate how topology can be realized in a biological system, given the many attractive properties of topology such as its robust response.
V.1.1 Long timescales: the circadian rhythm
A detailed case study of a topological mechanism in a concrete biological system is that of the KaiABC system for the circadian clock of cyanobacteria [89, 90]. As highlighted in Ref. 33, topological edge currents naturally as in Fig. 2(b) reproduce the kinetic ordering of KaiC phosphorylation cycles, although the detailed biophysical mechanisms remained abstract in this work. These missing mechanisms were made concrete (see Fig. 4) in Ref. 40, which also characterized the performance and efficiency of the resulting clock. In particular, it was shown that the key biochemical mechanisms [91, 92, 93, 94] interact via a separation of time-scales to produce the resulting edge currents, and provide the necessary and sufficient conditions for the observed oscillatory cycle [40]. An analysis of the parameter space showed how a variation of the transition rates (e.g. by changing ATP concentration [95] or using a mutant [96]) affects the coherence and dissipation of the oscillation, by tuning the system into and out of the topological transition. This yields key insights into an important regulatory system that had until now required rather complicated models, especially to reproduce the observed kinetic ordering of T and S phosphorylation [97, 98, 99].
On the theoretical level, Ref. 40 characterized the coherence of the resulting cycle, showing that it satisfies theoretical bounds [100] for the most coherent oscillator, equivalent to that of a unicyclic network, without the fine-tuning needed for a unicycle. The topological model shows high coherence compared to other available models, while producing the global day-night cycle with unusually few free parameters. In addition, the coherence and energetic cost of the oscillation were explored using tools from non-equilibrium stochastic thermodynamics, to reveal an efficient regime where coherence increases while cost simultaneously decreases. Lastly, a new indicator of oscillation coherence from spectral gaps in band theory was introduced, to study the saturation of this model on global thermodynamic bounds.
Overall, a topological model is able to explain long-standing puzzles in biology such as how dimensional reduction is achieved in a robust and flexible manner to produce emergent robust oscillations. More generally, it provides design principles for targeted dynamics in synthetic biological systems and in the engineering of reconfigurable materials, e.g., through dissipative self-assembly.
V.1.2 Sensory adaptation
In Ref. 35, the authors used their 1D topological model [Fig. 2(a)] to describe sensory adaption, in particular how E. coli senses chemical concentrations to navigate in response to chemical gradients [101, 102]. Here, the activity of the bacterial sensing complex is modified by its methylation level according to a sigmoid function. More specifically, there is low activity at low methylation, high activity at high methylation, and a narrow region in between where activity changes fast. The sensing complex tends to have a methylation level where activity changes the fastest, which ensures maximal sensitivity to chemoattractant.
The authors propose that their network, with a region of low methylation and activity and a region of high methylation and activity, can be described by a topological model with two different topological phases. The steady state is localized at the domain wall between the two phases. This ensures that the system is localized in the region where activity changes the fastest, allowing for rapid adaptation to environmental changes, while the topological origin for this mechanism makes the system robust against random perturbations.
V.1.3 Dynamical instability of microtubules
Using a 2D Kagome model with edge-localized currents [33], it was shown that similar dynamics to the growth and shrinkage of microtubules can arise from topologically-protected chiral edge currents. Specifically, the topological model predicts stable phases of growth followed by sporadic phases of shrinkage using only three parameters. While further work is needed to probe the specific biophysical features of the system, the model is able to successfully capture the wide range of catastrophe lengths and merits further investigation.
V.2 Synthetic microswimmers
The 2D topological model such as the one in Fig. 2(b) provides a novel design for low-Reynolds number microswimmers [54, 55], as depicted in Fig. 5 for the example of a stochastic three-sphere swimmer [64, 65]. By modeling the extension of the left and right arms of the swimmer as the two dimensions of our model, the topological edge state will naturally enclose the largest possible area in shape space, which is directly proportional to the self-propulsion velocity of the microswimmer [55, 64]. Notably, if either of the swimmer arms encounters a malfunction that e.g. limits its extension, the edge current would simply move around this obstacle to maintain the largest available phase space. Such models can be useful for the design of synthetic microswimmers or macroscopic robots in complex environments [103] that exhibit flexible behavior.
VI New theoretical tools and insights
The discovery of topological phases in stochastic systems has prompted the development of new theoretical tools to analyze them. This has revealed interesting new properties that are distinct from quantum and other topological systems, opening up new fields for investigation and study.
VI.1 Necessity of non-Hermiticity and out-of-equilibriumness
While there is strong motivation to realize topological states in stochastic systems as described above, general principles that govern stochastic topological systems remain lacking, together with an understanding of how stochastic topology generates edge responses necessary for robust function. Recent theoretical work has addressed this question and demonstrated that the only way for stochastic systems to exhibit a localized response is to have a non-Hermitian transition matrix , see Fig. 6 [73]. Note that Hermitian transition matrices always describe equilibrium systems with uniform steady states, whereas non-Hermitian ones can (but not necessarily do) describe out-of-equilibrium ones. This is in line with other phenomena that have no counterparts in equilibrium systems [38, 104, 16].
Interestingly, this non-Hermiticity condition is in sharp contrast with quantum systems, where Hermitian topological invariants are known to cause edge responses. This work establishes a strict condition for stochastic topological localization. Further, Ref. 73 showed that the edge response is markedly different in stochastic systems compared to quantum systems. They also demonstrate a novel mechanism by which non-Hermiticity engenders the steady-state current on the edge [73]. These findings reveal surprising differences between quantum and stochastic edge responses despite both systems sharing the same topological invariant.
VI.2 Topological invariants and relaxation time
The results above suggest the need for novel topological invariants designed for stochastic systems and their steady states, that signal more precisely the emergence of edge states. Such new tools have been developed for 1D stochastic systems [74, 75]. Specifically, the authors introduced a winding number using a scale-transformed transition-rate matrix in periodic boundary conditions, which predicts the existence of a spectral gap between the steady state and other states [74]. Further, this gap characterizes the relaxation time of the system towards the steady state: when the system is in the topological phase, the relaxation time scales with system size . In contrast, the trivial phase is expected to be gapless and have a relaxation time that scales as . Subsequently, the authors extended these results from ergodic systems also to non-ergodic ones, and developed a bulk-boundary correspondence between the winding number under periodic boundary conditions and the number of localized steady states in semi-infinite boundary conditions [75].
VI.3 Transfer matrix approach
The key quantities that define a topological state with chiral edge currents in a stochastic systems are the steady state accumulation of probability at the edges of the system, and the steady state edge current, i.e. how fast and in which direction this probability current flows along the edges of the system at steady state, see Fig. 2(b). In Ref. 33, a novel method to obtain these quantities was proposed, based on a transfer matrix approach. The key to the approach is to relate the steady state probabilities of neighboring cells to each other. In particular, for the model in Fig. 2(b), one can obtain a relation , where is a four-component vector containing the probabilities at the four sites of the unit cell that is a distance away from the edge of the system, and is a transfer matrix, independent of . The eigenvalues and eigenvectors of then contain all the necessary information about the edge states. In particular, was found to have two eigenvalues equal to one, corresponding to a uniform steady state of the bulk of the system, and two eigenvalues and , corresponding to edge modes located at opposite edges of the system. From the eigenvector associated to , the probability accumulation and edge current could then be explicitly calculated [33]. Additionally, corresponds to the decay length of the edge mode, which is damped by a factor in a unit cell a distance away from the edge.
VI.4 Topological classification using symmetries
The classification of topological phases has its origins in results from random matrix theory [8, 9, 7]. This is based on the study of modes that remain gapless under arbitrary perturbations of the Hamiltonian that preserve the characteristic symmetries, including disorder. It was found that the results depend only on the system’s symmetry, and Altland and Zirnbauer identified ten groups in total, an exhaustive list [105, 106, 107]. Specifically, they determined which surface modes can completely evade Anderson localization from random impurities, depending on the system dimension and symmetry class [8, 9].
The ten Altland-Zirnbauer groups form a classification of Hermitian matrices and their possible topological phases. Many of the early proposals for topological phases in discrete classical systems indeed map their systems into Hermitian Hamiltonians and use the topological properties associated to the relevant symmetry class and dimension [34, 36, 37, 70], employing the bulk-boundary correspondence to predict the appearance of edge states. However, stochastic systems governed by a master equation require an explicitly non-Hermitian description to exhibit surface states, when directly mapped to the Schrödinger equation [73]; see Fig. 6. There exist non-Hermitian features with no Hermitian counterpart, including the skin effect and an enhanced boundary sensitivity to random perturbations [71]. While some topological classifications of non-Hermitian Hamiltonians have been recently proposed [72], the relation of non-Hermitian topology to stochastic systems and the relevant symmetries for biological systems are questions that remain open.
VI.5 Thermodynamic properties and interactions
An interesting alternative approach using thermodynamic properties and the stoichiometric matrix associated to the transition matrix has been found, that can also predict the topological transition [108]. Separately, analysis of the entropy production in a topological model has revealed a regime of high efficiency, in which there is less dissipation despite greater coherence [40].
It would also be potentially useful to consider interactions between several biochemical or ecological systems. Inspired by biological systems under resource constraints, Ref. 33 considered coupled systems with shared boundaries, resulting in (anti)synchronized edge states. Going forward, the effects of different interactions and coupling, which effectively increases the dimension of the system, appears as a promising area for future investigation.
VII Outlook
We have highlighted several open theoretical directions in the previous section. Meanwhile, work remains to be done in the application and verification of existing topological models to realistic settings of biological relevance, as well as in the development of new models that exhibit yet unseen features that are exclusive to stochastic systems and have no quantum counterpart. It would also be useful to analyze the robustness of proposed models to forms of disorder, heterogeneity, and other constraints specified by real-world problems.
More broadly, this emerging field renders the powerful tools of topology and its robust dimensional reduction despite random perturbations, more accessible. Going forward, this research can lead to design prescriptions for targeted responses in synthetic biology, active matter, ecology, and robotics. This provides tools for building systems with stable function in the heterogeneous, non-equilibrium regimes which are the norm in everyday life.
Acknowledgements.
We thank Aleksandra Nelson for helpful comments. E.T. acknowledges support from the NSF Center for Theoretical Biological Physics (PHY-2019745) and the NSF CAREER Award (DMR-2238667).References
- Ross and Arkin [2009] J. Ross and A. P. Arkin, Complex systems: From chemistry to systems biology (2009).
- Ashkenasy et al. [2017] G. Ashkenasy, T. M. Hermans, S. Otto, and A. F. Taylor, Systems chemistry (2017).
- Dewyer et al. [2018] A. L. Dewyer, A. J. Argüelles, and P. M. Zimmerman, Methods for exploring reaction space in molecular systems, WIREs Computational Molecular Science 8, e1354 (2018).
- Davies [2020] P. Davies, Does new physics lurk inside living matter?, Phys. Today 73, 8 (2020).
- Winfree [1980] A. T. Winfree, The Geometry of Biological Time, Biomathematics (Berlin) (Springer Verlag, 1980).
- Gao and Ganguli [2015] P. Gao and S. Ganguli, On simplicity and complexity in the brave new world of large-scale neuroscience (2015).
- Mehta [1967] M. L. Mehta, Random matrices and the statistical theory of energy levels (Academic Press, 1967) p. 259.
- Schnyder et al. [2008] A. P. Schnyder, S. Ryu, A. Furusaki, and A. W. Ludwig, Classification of topological insulators and superconductors in three spatial dimensions, Physical Review B 78, 195125 (2008).
- Ryu et al. [2010] S. Ryu, A. P. Schnyder, A. Furusaki, and A. W. Ludwig, Topological insulators and superconductors: tenfold way and dimensional hierarchy, New Journal of Physics 12, 065010 (2010).
- Moore [2010] J. E. Moore, The birth of topological insulators (2010).
- Hasan and Kane [2010] M. Z. Hasan and C. L. Kane, Colloquium: Topological insulators, Reviews of Modern Physics 82, 3045 (2010).
- Chiu et al. [2016] C. K. Chiu, J. C. Teo, A. P. Schnyder, and S. Ryu, Classification of topological quantum matter with symmetries, Reviews of Modern Physics 88, 035005 (2016).
- Stormer et al. [1999] H. L. Stormer, D. C. Tsui, and A. C. Gossard, The fractional quantum hall effect, Reviews of Modern Physics 71, S298 (1999).
- Darcy [2021] I. Darcy, DNA Knots and Links, in Encyclopedia of Knot Theory (Chapman and Hall/CRC, 2021) pp. 901–910.
- Ardaševa and Doostmohammadi [2022] A. Ardaševa and A. Doostmohammadi, Topological defects in biological matter, Nature Reviews Physics 4, 354 (2022).
- Shankar et al. [2022] S. Shankar, A. Souslov, M. J. Bowick, M. C. Marchetti, and V. Vitelli, Topological active matter, Nature Reviews Physics 2022 4:6 4, 380 (2022).
- Tang and Fu [2014] E. Tang and L. Fu, Strain-induced partially flat band, helical snake states and interface superconductivity in topological crystalline insulators, Nature Physics 10, 964 (2014).
- Weis and Klitzing [2011] J. Weis and K. V. Klitzing, Metrology and microscopic picture of the integer quantum Hall effect, in Philosophical Transactions of the Royal Society A: Mathematical, Physical and Engineering Sciences, Vol. 369 (Royal Society, 2011) pp. 3954–3974.
- Cohen et al. [2019] E. Cohen, H. Larocque, F. Bouchard, F. Nejadsattari, Y. Gefen, and E. Karimi, Geometric phase from aharonov–bohm to pancharatnam–berry and beyond, Nat. Rev. Phys. 1, 437 (2019).
- Tang et al. [2011] E. Tang, J. W. Mei, and X. G. Wen, High-temperature fractional quantum hall states, Physical Review Letters 106, 236802 (2011).
- Obana et al. [2019] D. Obana, F. Liu, and K. Wakabayashi, Topological edge states in the Su-Schrieffer-Heeger model, Phys. Rev. B 100, 75437 (2019).
- Lieu [2018] S. Lieu, Topological phases in the non-Hermitian Su-Schrieffer-Heeger model, Phys. Rev. B 97, 45106 (2018).
- Su et al. [1979] W. P. Su, J. R. Schrieffer, and A. J. Heeger, Solitons in Polyacetylene, Phys. Rev. Lett. 42, 1698 (1979).
- Mao and Lubensky [2018] X. Mao and T. C. Lubensky, Maxwell Lattices and Topological Mechanics, Annual Review of Condensed Matter Physics 9, 413 (2018).
- Pedro et al. [2019] R. P. Pedro, J. Paulose, A. Souslov, M. Dresselhaus, and V. Vitelli, Topological Protection Can Arise from Thermal Fluctuations and Interactions, Phys. Rev. Lett. 122, 118001 (2019).
- Kane and Lubensky [2014] C. L. Kane and T. C. Lubensky, Topological boundary modes in isostatic lattices, Nat. Phys. 10, 39 (2014).
- Xiao et al. [2014] M. Xiao, Z. Q. Zhang, and C. T. Chan, Surface Impedance and Bulk Band Geometric Phases in One-Dimensional Systems, Phys. Rev. X 4, 21017 (2014).
- Benalcazar et al. [2017] W. A. Benalcazar, B. A. Bernevig, and T. L. Hughes, Quantized electric multipole insulators, Science 357, 61 (2017).
- Yuce and Ramezani [2019] C. Yuce and H. Ramezani, Topological states in a non-Hermitian two-dimensional Su-Schrieffer-Heeger model, Phys. Rev. A 100, 32102 (2019).
- Pocock et al. [2018] S. R. Pocock, X. Xiao, P. A. Huidobro, and V. Giannini, Topological Plasmonic Chain with Retardation and Radiative Effects, ACS Photonics 5, 2271 (2018).
- Ni et al. [2017] X. Ni, M. A. Gorlach, A. Alu, and A. B. Khanikaev, Topological edge states in acoustic kagome lattices, New J. Phys. 19, 055002 (2017).
- Liu et al. [2017] Y. Liu, C.-S. Lian, Y. Li, Y. Xu, and W. Duan, Pseudospins and Topological Effects of Phonons in a Kekulé Lattice, Phys. Rev. Lett. 119, 255901 (2017).
- Tang et al. [2021] E. Tang, J. Agudo-Canalejo, and R. Golestanian, Topology Protects Chiral Edge Currents in Stochastic Systems, Phys. Rev. X 11, 31015 (2021).
- Dasbiswas et al. [2018] K. Dasbiswas, K. K. Mandadapu, and S. Vaikuntanathan, Topological localization in out-of-equilibrium dissipative systems, Proc. Natl. Acad. Sci. U.S.A. 115, E9031 (2018).
- Murugan and Vaikuntanathan [2017] A. Murugan and S. Vaikuntanathan, Topologically protected modes in non-equilibrium stochastic systems, Nat. Commun. 8, 13881 (2017).
- Knebel et al. [2020] J. Knebel, P. M. Geiger, and E. Frey, Topological Phase Transition in Coupled Rock-Paper-Scissors Cycles, Phys. Rev. Lett. 125, 258301 (2020).
- Yoshida et al. [2021] T. Yoshida, T. Mizoguchi, and Y. Hatsugai, Chiral edge modes in evolutionary game theory: A kagome network of rock-paper-scissors cycles, Physical Review E 104, 025003 (2021).
- Hopfield [1974] J. J. Hopfield, Kinetic proofreading: a new mechanism for reducing errors in biosynthetic processes requiring high specificity, Proceedings of the National Academy of Sciences 71, 4135 (1974).
- Samoilov et al. [2005] M. Samoilov, S. Plyasunov, and A. P. Arkin, Stochastic amplification and signaling in enzymatic futile cycles through noise-induced bistability with oscillations, Proceedings of the National Academy of Sciences 102, 2310 (2005).
- Zheng and Tang [2023] C. Zheng and E. Tang, A topological mechanism for robust and efficient global oscillations in biological networks, arXiv:2302.11503 (2023).
- Nayak et al. [2008] C. Nayak, S. H. Simon, A. Stern, M. Freedman, and S. Das Sarma, Non-Abelian anyons and topological quantum computation, Reviews of Modern Physics 80, 1083 (2008).
- Thouless et al. [1982] D. J. Thouless, M. Kohmoto, M. P. Nightingale, and M. Den Nijs, Quantized hall conductance in a two-Dimensional periodic potential, Physical Review Letters 49, 405 (1982).
- Wang et al. [2014] C. Wang, A. C. Potter, and T. Senthil, Classification of interacting electronic topological insulators in three dimensions, Science 343, 629 (2014).
- Tang and Wen [2012] E. Tang and X. G. Wen, Interacting one-dimensional fermionic symmetry-protected topological phases, Physical Review Letters 109, 096403 (2012).
- DeCamp et al. [2015] S. J. DeCamp, G. S. Redner, A. Baskaran, M. F. Hagan, and Z. Dogic, Orientational order of motile defects in active nematics, Nature Materials 14, 1110 (2015).
- Keber et al. [2014] F. C. Keber, E. Loiseau, T. Sanchez, S. J. DeCamp, L. Giomi, M. J. Bowick, M. C. Marchetti, Z. Dogic, and A. R. Bausch, Topology and dynamics of active nematic vesicles, Science 345, 1135 (2014).
- Saw et al. [2017] T. B. Saw, A. Doostmohammadi, V. Nier, L. Kocgozlu, S. Thampi, Y. Toyama, P. Marcq, C. T. Lim, J. M. Yeomans, and B. Ladoux, Topological defects in epithelia govern cell death and extrusion, Nature 544, 212 (2017).
- Nicolau et al. [2011] M. Nicolau, A. J. Levine, and G. Carlsson, Topology based data analysis identifies a subgroup of breast cancers with a unique mutational profile and excellent survival, Proceedings of the National Academy of Sciences of the United States of America 108, 7265 (2011).
- Reimann et al. [2017] M. W. Reimann, M. Nolte, M. Scolamiero, K. Turner, R. Perin, G. Chindemi, P. Dłotko, R. Levi, K. Hess, and H. Markram, Cliques of neurons bound into cavities provide a missing link between structure and function, Frontiers in Computational Neuroscience 11, 10.3389/fncom.2017.00048 (2017).
- Taylor et al. [2015] D. Taylor, F. Klimm, H. A. Harrington, M. Kramár, K. Mischaikow, M. A. Porter, and P. J. Mucha, Topological data analysis of contagion maps for examining spreading processes on networks, Nature Communications 6, 1 (2015).
- Giusti et al. [2015] C. Giusti, E. Pastalkova, C. Curto, and V. Itskov, Clique topology reveals intrinsic geometric structure in neural correlations, Proceedings of the National Academy of Sciences of the United States of America 112, 13455 (2015).
- Giusti et al. [2016] C. Giusti, R. Ghrist, and D. S. Bassett, Two’s company, three (or more) is a simplex: Algebraic-topological tools for understanding higher-order structure in neural data (2016).
- Rocks et al. [2021] J. W. Rocks, A. J. Liu, and E. Katifori, Hidden Topological Structure of Flow Network Functionality, Physical Review Letters 126, 028102 (2021).
- Purcell [1977] E. M. Purcell, Life at low reynolds number, Am. J. Phys. 45, 3 (1977).
- Shapere and Wilczek [1987] A. Shapere and F. Wilczek, Self-propulsion at low reynolds number, Physical Review Letters 58, 2051 (1987).
- Shapere and Wilczek [1989] A. Shapere and F. Wilczek, Gauge kinematics of deformable bodies, American Journal of Physics 57, 514 (1989).
- Sinitsyn and Nemenman [2007] N. A. Sinitsyn and I. Nemenman, The Berry phase and the pump flux in stochastic chemical kinetics, EPL 77, 58001 (2007).
- May and Leonard [1975] R. M. May and W. J. Leonard, Nonlinear aspects of competition between three species, SIAM journal on applied mathematics 29, 243 (1975).
- Szolnoki et al. [2014] A. Szolnoki, M. Mobilia, L.-L. Jiang, B. Szczesny, A. M. Rucklidge, and M. Perc, Cyclic dominance in evolutionary games: a review, Journal of the Royal Society Interface 11, 20140735 (2014).
- Schnakenberg [1976] J. Schnakenberg, Network theory of microscopic and macroscopic behavior of master equation systems, Reviews of Modern physics 48, 571 (1976).
- Chodera and Noé [2014] J. D. Chodera and F. Noé, Markov state models of biomolecular conformational dynamics, Current opinion in structural biology 25, 135 (2014).
- Lipowsky et al. [2006] R. Lipowsky, Y. Chai, S. Klumpp, S. Liepelt, and M. J. Müller, Molecular motor traffic: From biological nanomachines to macroscopic transport, Physica A: Statistical Mechanics and its Applications 372, 34 (2006).
- Kolomeisky and Fisher [2007] A. B. Kolomeisky and M. E. Fisher, Molecular motors: a theorist’s perspective, Annu. Rev. Phys. Chem. 58, 675 (2007).
- Golestanian and Ajdari [2008] R. Golestanian and A. Ajdari, Mechanical response of a small swimmer driven by conformational transitions, Physical review letters 100, 038101 (2008).
- Chatzittofi et al. [2024] M. Chatzittofi, J. Agudo-Canalejo, and R. Golestanian, Entropy production and thermodynamic inference for stochastic microswimmers, Physical Review Research 6, L022044 (2024).
- Qian and Ge [2021] H. Qian and H. Ge, Stochastic Chemical Reaction Systems in Biology, Lecture Notes on Mathematical Modelling in the Life Sciences (Springer International Publishing, Cham, 2021).
- Goel et al. [1971] N. S. Goel, S. C. Maitra, and E. W. Montroll, On the volterra and other nonlinear models of interacting populations, Reviews of modern physics 43, 231 (1971).
- Knebel et al. [2015] J. Knebel, M. F. Weber, T. Krüger, and E. Frey, Evolutionary games of condensates in coupled birth–death processes, Nature communications 6, 6977 (2015).
- Yoshida et al. [2022] T. Yoshida, T. Mizoguchi, and Y. Hatsugai, Non-hermitian topology in rock–paper–scissors games, Scientific Reports 12, 560 (2022).
- Umer and Gong [2022] M. Umer and J. Gong, Topologically protected dynamics in three-dimensional nonlinear antisymmetric lotka-volterra systems, Physical Review B 106, L241403 (2022).
- Okuma and Sato [2023] N. Okuma and M. Sato, Non-Hermitian Topological Phenomena: A Review, Annual Review of Condensed Matter Physics 14, 83 (2023).
- Kawabata et al. [2019] K. Kawabata, K. Shiozaki, M. Ueda, and M. Sato, Symmetry and Topology in Non-Hermitian Physics, Phys. Rev. X 9, 41015 (2019).
- Nelson and Tang [2023] A. Nelson and E. Tang, Non-reciprocity permits edge states and strong localization in stochastic topological systems (2023), arXiv:2310.16720 [cond-mat.stat-mech] .
- Sawada et al. [2024a] T. Sawada, K. Sone, R. Hamazaki, Y. Ashida, and T. Sagawa, Role of Topology in Relaxation of One-Dimensional Stochastic Processes, Physical Review Letters 132, 046602 (2024a).
- Sawada et al. [2024b] T. Sawada, K. Sone, K. Yokomizo, Y. Ashida, and T. Sagawa, Bulk-Boundary Correspondence in Ergodic and Nonergodic One-Dimensional Stochastic Processes, arXiv (2024b).
- Rigotti et al. [2013] M. Rigotti, O. Barak, M. R. Warden, X.-J. Wang, N. D. Daw, E. K. Miller, and S. Fusi, The importance of mixed selectivity in complex cognitive tasks, Nature 497, 585 (2013).
- Tang et al. [2019] E. Tang, M. G. Mattar, C. Giusti, D. M. Lydon-Staley, S. L. Thompson-Schill, and D. S. Bassett, Effective learning is accompanied by high-dimensional and efficient representations of neural activity, Nature neuroscience 22, 1000 (2019).
- Stephens et al. [2008] G. J. Stephens, B. Johnson-Kerner, W. Bialek, and W. S. Ryu, Dimensionality and dynamics in the behavior of C. elegans, PLoS computational biology 4, e1000028 (2008).
- Chaudhuri and Fiete [2016] R. Chaudhuri and I. Fiete, Computational principles of memory, Nature neuroscience 19, 394 (2016).
- Hopfield [1982] J. J. Hopfield, Neural networks and physical systems with emergent collective computational abilities., Proceedings of the national academy of sciences 79, 2554 (1982).
- Liao and Rust [2021] Y. Liao and M. J. Rust, The circadian clock ensures successful dna replication in cyanobacteria, Proceedings of the National Academy of Sciences 118, e2022516118 (2021).
- Puszynska and O’Shea [2017] A. M. Puszynska and E. K. O’Shea, Switching of metabolic programs in response to light availability is an essential function of the cyanobacterial circadian output pathway, elife 6, e23210 (2017).
- Wolynes [2010] P. G. Wolynes, Gene Regulation: Single-Molecule Chemical Physics in a Natural Context, Springer Series in Chemical Physics 96, 553 (2010).
- Brettschneider et al. [2010] C. Brettschneider, R. J. Rose, S. Hertel, I. M. Axmann, A. J. Heck, and M. Kollmann, A sequestration feedback determines dynamics and temperature entrainment of the KaiABC circadian clock., Molecular systems biology 6, 389 (2010).
- Kageyama et al. [2006] H. Kageyama, T. Nishiwaki, M. Nakajima, H. Iwasaki, T. Oyama, and T. Kondo, Cyanobacterial Circadian Pacemaker: Kai Protein Complex Dynamics in the KaiC Phosphorylation Cycle In Vitro, Molecular Cell 23, 161 (2006).
- Lin et al. [2014] J. Lin, J. Chew, U. Chockanathan, and M. J. Rust, Mixtures of opposing phosphorylations within hexamers precisely time feedback in the cyanobacterial circadian clock, Proceedings of the National Academy of Sciences of the United States of America 111, E3937 (2014).
- Xu et al. [2003] Y. Xu, T. Mori, and C. H. Johnson, Cyanobacterial circadian clockwork: roles of KaiA, KaiB and the kaiBC promoter in regulating KaiC, The EMBO Journal 22, 2117 (2003).
- Rust et al. [2007] M. J. Rust, J. S. Markson, W. S. Lane, D. S. Fisher, and E. K. O’Shea, Ordered phosphorylation governs oscillation of a three-protein circadian clock, Science 318, 809 (2007).
- Ishiura et al. [1998] M. Ishiura, S. Kutsuna, S. Aoki, H. Iwasaki, C. R. Andersson, A. Tanabe, S. S. Golden, C. H. Johnson, and T. Kondo, Expression of a gene cluster kaiABC as a circadian feedback process in cyanobacteria, Science 281, 1519 (1998).
- Nakajima et al. [2005] M. Nakajima, K. Imai, H. Ito, T. Nishiwaki, Y. Murayama, H. Iwasaki, T. Oyama, and T. Kondo, Reconstitution of circadian oscillation of cyanobacterial KaiC phosphorylation in vitro, science 308, 414 (2005).
- Kim et al. [2008] Y.-I. Kim, G. Dong, C. W. Carruthers Jr, S. S. Golden, and A. LiWang, The day/night switch in KaiC, a central oscillator component of the circadian clock of cyanobacteria, Proceedings of the National Academy of Sciences 105, 12825 (2008).
- Chang et al. [2011] Y.-G. Chang, N.-W. Kuo, R. Tseng, and A. LiWang, Flexibility of the C-terminal, or CII, ring of KaiC governs the rhythm of the circadian clock of cyanobacteria, Proceedings of the National Academy of Sciences 108, 14431 (2011).
- Tseng et al. [2014] R. Tseng, Y.-G. Chang, I. Bravo, R. Latham, A. Chaudhary, N.-W. Kuo, and A. LiWang, Cooperative KaiA–KaiB–KaiC interactions affect KaiB/SasA competition in the circadian clock of cyanobacteria, Journal of molecular biology 426, 389 (2014).
- Hong et al. [2018] L. Hong, B. P. Vani, E. H. Thiede, M. J. Rust, and A. R. Dinner, Molecular dynamics simulations of nucleotide release from the circadian clock protein kaic reveal atomic-resolution functional insights, Proceedings of the National Academy of Sciences 115, E11475 (2018).
- Phong et al. [2013] C. Phong, J. S. Markson, C. M. Wilhoite, and M. J. Rust, Robust and tunable circadian rhythms from differentially sensitive catalytic domains, Proceedings of the National Academy of Sciences 110, 1124 (2013).
- Qin et al. [2010] X. Qin, M. Byrne, T. Mori, P. Zou, D. R. Williams, H. Mchaourab, and C. H. Johnson, Intermolecular associations determine the dynamics of the circadian KaiABC oscillator, Proceedings of the National Academy of Sciences 107, 14805 (2010).
- van Zon et al. [2007] J. S. van Zon, D. K. Lubensky, P. R. Altena, and P. R. ten Wolde, An allosteric model of circadian KaiC phosphorylation, Proceedings of the National Academy of Sciences 104, 7420 (2007).
- Paijmans et al. [2017] J. Paijmans, D. K. Lubensky, and P. R. Ten Wolde, A thermodynamically consistent model of the post-translational Kai circadian clock, PLoS computational biology 13, e1005415 (2017).
- Smolen et al. [2001] P. Smolen, D. A. Baxter, and J. H. Byrne, Modeling circadian oscillations with interlocking positive and negative feedback loops, Journal of Neuroscience 21, 6644 (2001).
- Barato and Seifert [2017] A. C. Barato and U. Seifert, Coherence of biochemical oscillations is bounded by driving force and network topology, Phys. Rev. E 95, 062409 (2017).
- Barkai and Leibler [1997] N. Barkai and S. Leibler, Robustness in simple biochemical networks, Nature 387, 913 (1997).
- Tu [2013] Y. Tu, Quantitative Modeling of Bacterial Chemotaxis: Signal Amplification and Accurate Adaptation, Annual Review of Biophysics 42, 337 (2013).
- Hatton et al. [2013] R. L. Hatton, Y. Ding, H. Choset, and D. I. Goldman, Geometric visualization of self-propulsion in a complex medium, Physical Review Letters 110, 078101 (2013).
- Qian and Qian [2000] H. Qian and M. Qian, Pumped biochemical reactions, nonequilibrium circulation, and stochastic resonance, Phys. Rev. Lett. 84, 2271 (2000).
- Wigner [1955] E. P. Wigner, Characteristic Vectors of Bordered Matrices With Infinite Dimensions, The Annals of Mathematics 62, 548 (1955).
- Dyson [1962] F. J. Dyson, The Threefold Way. Algebraic Structure of Symmetry Groups and Ensembles in Quantum Mechanics, Journal of Mathematical Physics 3, 1199 (1962).
- Altland and Zirnbauer [1997] A. Altland and M. R. Zirnbauer, Nonstandard symmetry classes in mesoscopic normal-superconducting hybrid structures, Physical Review B 55, 1142 (1997).
- Mehta and Rocks [2022] P. Mehta and J. Rocks, Thermodynamic origins of topological protection in nonequilibrium stochastic systems, arXiv (2022).